OpenStudy (anonymous):

Geometry the triangle has an area of 40cm squared.Find the height H of the triangle.

OpenStudy (hba):

Assume base length = 2h +6 Area = 1/2 * base * height 40 = 1/2 * (2h+6) * h 40 = (h+3)h 40= h^2 +3h h^2 +3h -40 = 0 ( h +8)(h -5) = 0 so h = 5 or -8 But h > 0 so h = 5 Therefore h = 5 cm

OpenStudy (lgbasallote):

@hba how did base length become 2h + 6?

OpenStudy (unklerhaukus):

how can you assume the base length @hba??

OpenStudy (hba):

cant we assume

OpenStudy (lgbasallote):

in math we dont assume

OpenStudy (anonymous):

yes until you put "h" in it

OpenStudy (anonymous):

Thanks For Your Good answers :)

OpenStudy (lgbasallote):

@hba in the case of yahoo answers ( http://answers.yahoo.com/question/index?qid=20120212152005AALS3Tw ) there was a given value of length in this case where the length is not given you cannot assume it's that

OpenStudy (hba):

k

OpenStudy (lgbasallote):

however @KateÜü if your question is same to yahoo answers...then you can assume that

OpenStudy (hba):

thanks

OpenStudy (anonymous):

:)